OpenStudy (kj4uts):

Reflecting the graph y=sin x across the y-axis is the same as reflecting it across the x-axis. True or False

OpenStudy (anonymous):

FALSE

OpenStudy (jhannybean):

Draw out what the sine graph looks like first.

OpenStudy (anonymous):

false

OpenStudy (anonymous):

why don't we actually check it instead of guessing

OpenStudy (anonymous):

i will send two pictures one of \[-\sin(x)\] and the other of \[\sin(-x)\] and we can see if they are the same

OpenStudy (anonymous):

before we go ahead an say "false" maybe we can remember that sine is an odd function, so it is always the case that \[\sin(-x)=-\sin(x)\]